# A New Impossibility Result for Online Bipartite Matching Problems

### v1

@@ 8 Generalization to Non-Equibipartite Graphs
The following claim follows as a corollary from the analysis of the maximum matching for the case $u=1$. {corollaryrep} The maximum matching of the instance of Footnote 10 has size $u\cdot n-o(n)$ with probability $1-o(1)$; thus, the resulting graph admits a quasi-complete matching. {appendixproof} We only need to prove this for $u<1$. Let us add $n-\lceil u\cdot n\rceil$ new users: $\frac{n}{\Delta}$ users of degree 0 and $(1-u-\frac{1}{\Delta})n$ users of degree $\Delta$ (rounded to be integers). With an argument similar to Section 6.2.1, for each $d\in\{2,3,\dots,\Delta\}$ the number of users of degree $d$ is concentrated around the expected value. Conditioning on this event that happens with probability $1-o(1)$, the number of edges is $\sum_{d=2}^{\Delta}\frac{n}{d-1}\pm o(n^{\nicefrac{{4}}{{5}}})=(H_{\Delta-1}\pm o(n^{\nicefrac{{-1}}{{5}}}))n$. Therefore, with probability $1-o(1)$, the degrees of the ads follow a Poisson distribution of parameter $H_{\Delta-1}$ as in Section 6.2.1. Conditioning on this event, we obtain the same configuration model as the one of Section 6.2.1, therefore the resulting graph has a matching of size at least $(1-\frac{1}{\Delta}-o(1))n$, and thus, the original $D_{u}$ instance has a matching of size at least $(1-\frac{1}{\Delta}-1+u+\frac{1}{\Delta}-o(1))n=(u-o(1))n$, with probability $1-o(1)$.
@@
Thus, the resulting graph is the configuration model described in Section 7 with $\Delta=i$ and therefore, for a small enough constant $\epsilon$ with respect to $i$, with probability $1-o(1)$ the maximum matching has size at most $(1-\frac{1}{i}+\epsilon-\psi_{i}(\epsilon)+o(1))n$, for $\psi_{i}(\epsilon)>0$. Thus, in the instance given by $P$, with probability $1-o(1)$, the number of users of degree $\leq i$ that can be matched is at most $(1-\frac{1}{i}+\epsilon-\psi_{i}(\epsilon)+o(1))n+(\rho u-\epsilon+o(1))n=(\Pr[P\leq i]-\frac{\psi_{i}(\epsilon)}{u})u\cdot n+o(n)$. Note that $0<\psi_{i}(\epsilon)\leq\epsilon<u\rho$ and $\nicefrac{{\psi_{i}(\epsilon)}}{{u}}>0$, which concludes the proof. We now move on to studying the performance of online algorithms on the generalized instance of Footnote 10. We start with a technical lemma that relates the expected number of users to capture a new ad with distribution $D_{u}$, and with distribution $D$. This Lemma will be used to bound the number of users needed to capture a new ad with distribution $D_{u}$, in terms of the number of users needed with $D$ — we need this Lemma because the integral required to determine the number of users to capture an ad with $D_{u}$ is not easy to solve, while the one with $D$ has already been solved in the proof of Section 6.1. {lemmarep} Fix $u\in\left(\frac{1}{2},1\right)$ and let $\Delta=\left\lceil\frac{1}{1-u}\right\rceil$, $c_{i}=\frac{1}{u\cdot i\cdot(i-1)}$ for $i\in\{2,\ldots,\Delta-1\}$ and $c_{\Delta}=1-\sum_{i=2}^{\Delta-1}\frac{1}{u\cdot i\cdot(i-1)}=1-\frac{\Delta-2}{u\cdot(\Delta-1)}$. [^11] Observe that $c_{i}=\Pr[D_{u}=i]\;\forall i\in\{2,\ldots,\Delta\}$, where $D_{u}$ is the distribution of Definition 10. Moreover, let $x\in\left[0,u\right)$. Then,

### v2

@@ 8 Generalization to Non-Equibipartite Graphs
The following claim is a corollary of the analysis of the maximum matching for the case $u=1$. {corollaryrep} The maximum matching of the instance of Footnote 10 has size $u\cdot n-o(n)$ with probability $1-o(1)$; thus, the resulting graph admits a quasi-complete matching. {appendixproof} We only need to prove this for $u<1$. Let us add $n-\lceil u\cdot n\rceil$ new users: $\frac{n}{\Delta}$ users of degree 0 and $(1-u-\frac{1}{\Delta})n$ users of degree $\Delta$ (rounded to be integers). With an argument similar to Section 6.2.1, for each $d\in\{2,3,\dots,\Delta\}$ the number of users of degree $d$ is concentrated around the expected value. Conditioning on this event that happens with probability $1-o(1)$, the number of edges is $\sum_{d=2}^{\Delta}\frac{n}{d-1}\pm o(n^{\nicefrac{{4}}{{5}}})=(H_{\Delta-1}\pm o(n^{\nicefrac{{-1}}{{5}}}))n$. Therefore, with probability $1-o(1)$, the degrees of the ads follow a Poisson distribution of parameter $H_{\Delta-1}$ as in Section 6.2.1. Conditioning on this event, we obtain the same configuration model as the one of Section 6.2.1, therefore the resulting graph has a matching of size at least $(1-\frac{1}{\Delta}-o(1))n$, and thus, the original $D_{u}$ instance has a matching of size at least $(1-\frac{1}{\Delta}-1+u+\frac{1}{\Delta}-o(1))n=(u-o(1))n$, with probability $1-o(1)$.
@@
Thus, the resulting graph is the configuration model described in Section 7 with $\Delta=i$ and therefore, for a small enough constant $\epsilon$ with respect to $i$, with probability $1-o(1)$ the maximum matching has size at most $(1-\frac{1}{i}+\epsilon-\psi_{i}(\epsilon)+o(1))n$, for $\psi_{i}(\epsilon)>0$. Thus, in the instance given by $P$, with probability $1-o(1)$, the number of users of degree $\leq i$ that can be matched is at most $(1-\frac{1}{i}+\epsilon-\psi_{i}(\epsilon)+o(1))n+(\rho u-\epsilon+o(1))n=(\Pr[P\leq i]-\frac{\psi_{i}(\epsilon)}{u})u\cdot n+o(n)$. Note that $0<\psi_{i}(\epsilon)\leq\epsilon<u\rho$ and $\nicefrac{{\psi_{i}(\epsilon)}}{{u}}>0$, which concludes the proof. We now move on to studying the performance of online algorithms on the generalized instance of Footnote 10. We start with a technical lemma that relates the expected number of users to capture a new ad with distribution $D_{u}$, and with distribution $D$. This Lemma will be used to bound the number of users needed to capture a new ad with distribution $D_{u}$, in terms of the number of users needed with $D$ — we need this Lemma because the integral required to determine the number of users to capture an ad with $D_{u}$ is not easy to solve, while the one for $D$ has already been solved in the proof of Section 6.1. {lemmarep} Fix $u\in\left(\frac{1}{2},1\right)$ and let $\Delta=\left\lceil\frac{1}{1-u}\right\rceil$, $c_{i}=\frac{1}{u\cdot i\cdot(i-1)}$ for $i\in\{2,\ldots,\Delta-1\}$ and $c_{\Delta}=1-\sum_{i=2}^{\Delta-1}\frac{1}{u\cdot i\cdot(i-1)}=1-\frac{\Delta-2}{u\cdot(\Delta-1)}$. [^11] Observe that $c_{i}=\Pr[D_{u}=i]\;\forall i\in\{2,\ldots,\Delta\}$, where $D_{u}$ is the distribution of Definition 10. Moreover, let $x\in\left[0,u\right)$. Then,